# Planar near-field measurements of specular and diffuse reflection of millimeter-wave absorbers

### v1

@@ 4.2 Angular responses and reflectance of the absorbers
We also compared the measured specular reflectance discussed in Sec. 2.1 with conventional horn-to-horn measurements. We performed the horn-to-horn measurements using the same experimental setup in Fig. 1[Sec. 3.1] by replacing the probe with the lens horn. The two lens horns for the transmitter and receiver are identical. The results demonstrate that the planar near-field measurement and the horn-to-horn reflection measurement agree within a few dB of accuracy and that the near-field measurements are less affected by standing waves (Fig. 9).
@@ 4.3 Electric field measurements perpendicular to the co-polarization of the lens horn
Using the planar near-field measurement method, we also measured the electric field perpendicular to the co-polarization of the lens horn. Figure 10 shows the angular responses of the electric field perpendicular to the co-polarization of the lens horn at 90 GHz for AoI = 30∘. The patterns of AN-72, LS-22, and TK-S measured with the S- and P-polarized lens horn (labeled as “sp” and “ps” respectively) are similar to those of the reference aluminum plate (“sp” and “ps” in Fig. 3) with attenuated levels, showing a four-divided pattern. This indicates that the patterns are determined mainly by the cross-polarization of the lens horn. On the other hand, the patterns of HR-10 and TK-L do not resemble that of the reference aluminum plate.
Such trends are also apparent in total reflectance defined by Eq. (3), as summarized in Fig. 11. For AN-72, LS-22, and TK-S, the total reflectance measured with the S- and P-polarized lens horn ( and respectively) are mostly consistent with and . Here, and indicate the cross-polarization components of the lens horn, which are obtained from the reference aluminum plate measurements as,
However, the total reflectance and of HR-10 and TK-L are dB larger than and . The measurements imply that HR-10 and TK-L may affect the polarization direction when reflecting the incident waves.
@@ 5 Conclusion
While the conventional horn-to-horn method measures the power coupling efficiency between the two horns, the present method allows the detailed near-field mapping of the wave reflected from the absorber. Thus, the latter can investigate the nature of the reflection more extensively. We have demonstrated in this paper that the diffuse-to-specular reflectance ratio is different among various materials. It is also a merit of the new method that the measurement is not sensitive to the alignment of the absorber in the setup optics. Practically, it will help us conduct similar reflection measurements in cryogenic conditions. The method is expected to contribute to the development of millimeter-wave absorbers by achieving accurate measurements, including diffuse reflections, and to make a better design against the stray light in CMB observation telescopes.
@@
Acknowledgments We are grateful to Luca Lamagna, Andrea Occhiuzzi, Haruaki Hirose and Takuro Fujino for their helpful discussions on the measurements. We are also grateful to Oshima Prototype Engineering Co. for fabricating the horn antennas used in the measurements.

### v2

@@ 4.4 Comparison with horn-to-horn method
We compared the measured specular reflectance discussed in Sec. 2.1 with conventional horn-to-horn measurements. We performed the horn-to-horn measurements using the same experimental setup described in Sec. 3.1 by replacing the probe with the lens horn. The two lens horns for the transmitter and receiver are identical. The results demonstrate that the planar near-field measurements and the horn-to-horn reflection measurements agree within a few dB and that the former are less affected by standing waves (Fig. 9). The measurements are consistent with the specular reflectance of TK-S reported in Ref. [9], measured with the horn-to-horn method at 90 – 110 GHz.
On the other hand, compared to the horn-to-horn method, the near-field method takes longer for each measurement. Also, near-field measurements require a wide scan plane and thus need a larger mechanical clearance between the transmitter and receiver, which limits the measurable angle of incidence.
@@ 4.5 Electric field measurements perpendicular to the co-polarization of the lens horn
Using the planar near-field measurement method, we also measured the electric field perpendicular to the co-polarization of the lens horn. Figure 10 shows the angular responses of the electric field perpendicular to the co-polarization of the lens horn at 90 GHz for AoI = 30∘. The patterns of AN-72, LS-22, and TK-S measured with the S- and P-polarized lens horn (labeled as “sp” and “ps” respectively) are similar to those of the reference aluminum plate (“sp” and “ps” in Fig. 3) with attenuated levels, showing quadrupoles. This indicates that the patterns are determined mainly by the cross-polarization of the lens horn. On the other hand, the patterns of HR-10 and TK-L do not resemble that of the reference aluminum plate.
Such trends are also apparent in total reflectance defined by Eq. (3), as summarized in Fig. 11. For AN-72, LS-22, and TK-S, the total reflectance measured with the S- and P-polarized lens horn ( and respectively) are primarily consistent with and . Here, and indicate the cross-polarization components of the lens horn, which are obtained from the reference aluminum plate measurements as,
In contrast, the total reflectance and of HR-10 and TK-L are dB larger than and . These measurements imply that HR-10 and TK-L may affect the polarization direction when reflecting the incident waves.
@@ 5 Conclusion
While the conventional horn-to-horn method measures the optical coupling efficiency between the two horns, the present method allows the detailed near-field mapping of the wave reflected from the absorber. Thus, the latter can investigate the nature of the reflection more extensively. We have demonstrated in this paper that the diffuse-to-specular reflectance ratio is different among various materials. It is also a merit of the new method that the measurement is not sensitive to the alignment of the absorber in the setup optics. Practically, it will help us conduct similar reflection measurements in cryogenic conditions. The method is expected to contribute to the development of millimeter-wave absorbers by achieving accurate measurements, including diffuse reflections, and to make a better design against the stray light in CMB observation telescopes.
@@
Acknowledgments We are grateful to Luca Lamagna, Andrea Occhiuzzi, Haruaki Hirose, Takuro Fujino, Rion Takahashi and Karen Sakamoto for their helpful discussions on the measurements. We are also grateful to Oshima Prototype Engineering Co. for fabricating the horn antennas used in the measurements.